# Can AI be enabled to dynamical downscaling? Training a Latent Diffusion Model to mimic km-scale COSMO-CLM downscaling of ERA5 over Italy

### v1

@@ Abstract
Downscaling techniques are one of the most prominent applications of Deep Learning (DL) in Earth System Modeling. A robust DL downscaling model can generate high-resolution fields from coarse-scale numerical model simulations, saving the timely and resourceful applications of regional/local models. Additionally, generative DL models have the potential to provide uncertainty information, by generating ensemble-like scenario pools, a task that is computationally prohibitive for traditional numerical simulations. In this study, we apply a Latent Diffusion Model (LDM) to downscale ERA5 data [1] over Italy up to a resolution of 2 km. The high-resolution target data consists of results from a high-resolution dynamical downscaling performed with COSMO-CLM [2, 3]. Our goal is to demonstrate that recent advancements in generative modeling enable DL-based models to deliver results comparable to those of numerical dynamical downscaling models, given the same input data (i.e., ERA5 data), preserving the realism of fine-scale features and flow characteristics. The training and testing database consists of hourly data from 2000 to 2020 ( 184000 timestamps). The target variables of this study are 2-m temperature and 10-m horizontal wind components. A selection of predictors from ERA5 is used as input to the LDM (e.g., 850hPa temperature, specific humidity, and wind), and a residual approach against a reference UNET is leveraged in applying the LDM. The performance of the generative LDM is compared with reference baselines of increasing complexity: quadratic interpolation of ERA5, a UNET, and a Generative Adversarial Network (GAN) built on the same reference UNET. Results highlight the improvements introduced by the LDM architecture and the residual approach over these baselines. The models are evaluated on a yearly test dataset, assessing the models’ performance through deterministic metrics (e.g. RMSE, BIAS, R2, …), spatial distribution of errors, and reconstruction of frequency and power spectra distributions.
Keywords Downscaling  $\cdot$ Deep learning  $\cdot$ Dynamical reanalyses  $\cdot$ Generative models
1 Introduction
@@
High-resolution near-surface meteorological fields such as 2-m temperature and 10-m wind speed are key targets for the weather and climate scientific and operational communities. Such high-resolution information is of essential importance for a wide variety of applications (e.g. available wind potential, predicted energy consumption, water availability, …), across diverse timescales, from weather forecasting (nowcasting, medium-range forecasting, and seasonal predictions) to climate projections. The hunger for high-resolution data is directly linked to and justified by the information that such data hold: extreme weather events and localized phenomena can typically be described by highly resolved fields only.
@@
Downscaling is a well-known approach that allows obtaining local high-resolution data (predictands) starting from low-resolution information (predictors) by applying suitable refinement techniques. The two most traditional approaches are dynamical downscaling and statistical downscaling ([4], [5], [6]), applied alternatively depending on the final goal of each specific application.
@@
Traditionally, high-resolution fields are achieved in operational weather forecasting by performing dynamical downscaling of lower-resolution simulations. Examples of this approach are all the Local Area Models (LAMs) run in every operational center worldwide, fed with global circulation models at a low resolution, and producing high-resolution fields for a localized area, typically nationwide (e.g. [7], [8], [9]). As for the climate community, this approach materializes, for example, in applications run within the WCRP Coordinated Regional Downscaling Experiment (CORDEX, [10]), performing dynamical downscaling of climate projections with Regional Climate Models (RCM) going from the $\sim$100-km resolution of Global Climate Models down to a $\sim$16-km resolution (e.g. [11]). The dynamical downscaling approach is well-established and provides physically and temporally consistent fields. However, it still has significant drawbacks due to the high resource demands required for its execution that limit its application, e.g. to deterministic runs instead of (or limited to small) ensemble runs.
@@
On the other hand, statistical downscaling uses coarse data from numerical simulations to infer data at high resolution by applying empirical relationships or transfer functions derived from a set of known predictors–predictands data pairs ([6]). Statistical downscaling methods have evolved through the years since the 1990s with increasingly greater levels of complexity and amount of data. Canonical examples of statistical downscaling methods are linear or multilinear regression methods (e.g. [12], [13]), analog ensemble downscaling (e.g. [14]) or quantile mapping ([15]).
@@
In recent years, the advent of machine-learning techniques introduced many other powerful methods. These approaches have the potential to outperform classical statistical models, introducing nonlinear components in the downscaling process and learning from the provided high-resolution data. Specifically, Convolutional Neural Networks (CNNs) are particularly well suited for handling spatially distributed data and for the super-resolution task, being able to capture complex, nonlinear mappings identifying crucial features, and have already been successfully applied to weather downscaling (e.g. [16, 17, 18]). Building on CNN frameworks, two Deep Learning approaches are currently the most promising for improving atmospheric downscaling: Generative Adversarial Networks (GANs, [19, 20]) and Diffusion Models ([21]), which both allow for a probabilistic approach to the problem. The potential and drawbacks of these approaches are widely reported in the following Section 2.
@@
In this work, we focus on and test a Latent Diffusion model that is novel for application in the atmospheric downscaling task. Its advantages should be twofold: the diffusion approach has a much more stable training than that of GAN models, still holding the ability to generate small-scale features and the potential for ensemble production, and shows superior results in applications to image processing compared to GANs ([22, 23]). In addition, the latent approach should be seen as incremental compared to pixel-space-based diffusion approaches as it provides a cheaper solution in terms of computational costs for both inference and training ([24]). This feature is intrinsically appealing to scale the downscaling task to wider (and longer) domains. Lastly, we use the high-resolution output from a numerical model dynamical downscaling simulation as our target-reference-truth. In doing so, we aim to determine whether a properly trained deep learning model can effectively emulate dynamical downscaling. If successful, this model could serve as a versatile dynamical downscaling emulator, much faster than the traditional numerical dynamical models, with a broad range of utmost significant applications.
@@ 2 Related work and Contribution
Building on CNN frameworks, two Deep Learning (DL) approaches are currently the most promising for improving atmospheric downscaling: Generative Adversarial Networks (GANs) and Diffusion Models, which both allow for the generation of small-scale features and a probabilistic approach to the problem. GANs have already shown promising results in downscaling different meteorological variables, in different regions [25, 26, 27] but pose relevant challenges (e.g., instabilities, mode collapses) during the training procedure ([28, 29]).
@@
On the other hand, Diffusion models introduce a relatively younger approach and have already been proven very effective in weather forecasting and nowcasting applications (e.g. [30, 31]). Diffusion models have yet to be widely tested and evaluated on the atmospheric downscaling task, but their characteristics and capabilities are undoubtedly promising for this application as shown for example in [32, 33].
@@
Building on these encouraging results, in this work, we approach the downscaling task with a Latent diffusion model, comparing it against some standard baselines and a GAN baseline. Specifically, we re-adapted the Latent Diffusion Cast (LDCast) model [30], recently developed for precipitation nowcasting. LDCast has shown superior performance in the generation of highly realistic precipitation forecast ensembles, and in the representation of uncertainty, compared to traditional Generative Adversarial Network (GAN)-based methods. Our resulting model for downscaling, similar to fully convolutional models, can be trained on examples of smaller spatial domains (patches) and used at the evaluation stage on domains of arbitrary sizes, making it suitable for the generation of high-resolution data covering wider domains. As also suggested in [33], we propose the application of the diffusion model with a residual approach, relying on a standard UNET architecture for capturing the bigger scales and training the latent diffusion model to generate the residual, small scales only.
@@
Our work differs from most of the above-mentioned works also in terms of the chosen pair of low-resolution and high-resolution data for the training. This choice highly influences the level of complexity that the DL downscaling model must achieve. Indeed, downscaling a coarsening of the high-resolution data (e.g. [27, 25, 34]) is a much easier task than downscaling modeled low-resolution data (e.g. short-term forecasts as in [26], seasonal predictions, or climate projections) to independent high-resolution data, either coming from observations or numerical model simulation. While the first exercise falls into a purely super-resolution task, the latter includes learning model biases and correcting them, including the identification of the inability of large-scale numerical models to detect local phenomena that cannot be resolved at their coarse resolution. In our work, we focus on reanalyses products and we train our models using a set of 14 ERA5 variables as low-resolution input and high-resolution data from a dynamical downscaling of ERA5 (run with the COSMO-CLM model) as target data (2-m temperature and 10-m wind speed components). This approach is similar to that followed, for example, by [35, 33]. In doing so, we intentionally force the model to learn to generate the effects of those local phenomena resolved by the dynamical numerical model, emulating its behavior.
@@ 3.1 Low- and High-resolution data
The goal of this experiment is to train DL models to mimic a dynamical downscaling performed with a convection-permitting Regional Climate Model (RCM). The target high-resolution data consists of the hourly Italian Very High-Resolution Reanalyses produced with COSMO5.0_CLM9 (VHR-REA CCLM) by dynamically downscaling ERA5 reanalyses [1] from its native resolution ( 25km) up to  2.2km over Italy ([2, 3]). Consistently with these target numerical simulations, the input low-resolution data fed to our DL models are ERA5 data.
@@ 3.2 Data alignment and preprocessing
ERA5 data have a resolution of 0.25° worldwide, which roughly corresponds to  22 km at the latitudes of the focus domain, while VHR-REA CCLM data have a native resolution of 0.02° (2.2 km). Data from both datasets were preprocessed to re-project, trim, and align the low and high-resolution fields. Specifically, the coordinate reference system (CRS) chosen for the experiment is ETRS89-LAEA Europe (Lambert Azimuthal Equal Area), also known in the EPSG Geodetic Parameter Dataset under the identifier EPSG:3035 ([36]), and the experiment grids align with the European Environmental Agency Reference grid (EEA Reference grid, [37]). ERA5 was reprojected and interpolated (with nearest-neighbor interpolation) on the EEA 16-km reference grid while VHR-REA CCLM was reprojected and interpolated on the EEA 2-km reference grid. The factor of the downscaling procedure is, therefore, 8x: over the target domain, low-resolution data consists of 72x86 16-km pixels images while high-resolution data consists of 576x672 2-km pixels images.
@@ 3.3 Experimental domain
The experiment target domain spans from 36°N to 48°N, and from 5°E to 20°E (Figure 1). This area corresponds to the region where VHR-REA CCLM data are available. The region includes a wide variety of topographically different sub-areas (mountainous areas such as the Alps and the Appennini, flat areas such as the Po Valley and coastal lines) which trigger local phenomena whose effects are challenging to identify for the DL downscaling models, as they are not present in the low-resolution data.
@@ 3.4 Target variables and predictors
The target variables of the study are (i) 2-m temperature and (ii) horizontal wind components at 10 m: different, dedicated models to each target variable have been trained. The input ERA5 low-resolution data are the target variables and additional fields used as dynamical predictors, to improve models’ performance. The choice of the set of input fields was based on previous literature (e.g. [18], [17], [26]) and on hardware constraints for the experiment. The selected fields used as predictor variables, for both target variables, are the following, corresponding to a total of 14 input channels to our networks:
@@
  • pressure mean sea level
@@
DEM data consists of the Copernicus Digital Elevation Model (DEM, [38]) interpolated from a resolution of 90 m to a resolution of 2 km. Land cover data were retrieved from the Copernicus Land Service, Global Land Cover data (GLC, [39]) interpolated from a resolution of 100 m to a resolution of 2 km. Given that land cover was utilized as a static variable in our analysis, we selected data from 2015: this year represents the earliest epoch available for the selected GLC dataset and falls approximately amid our experimental period of 2000-2020. Land cover class data have been converted to single-channel class masks for the DL models (totaling 16 channels). All static fields have been re-projected and aligned to the high-resolution 2-km EAA reference grid.
@@ 4 Methods
In this work, we test a deep generative Latent Diffusion Model (LDM) for the downscaling task, conditioned with low-resolution predictors and high-resolution static data. The implemented LDM is trained to predict the residual error between a previously trained reference UNET and the target variables, hence the model will be addressed as LDM_res (LDM_residual) hereafter. This residual approach has shown great performance in the application of pixel-space diffusion models [33] and is tested here in the latent diffusion context. The underlying idea is to exploit the great ability of a relatively simple network (a UNET) to properly capture the main, bigger-scale variability of the atmospheric high-resolution data and leverage the power of the generative diffusion model to only focus on the reconstruction of the smaller-scale, locally-driven variability of the fields. Figure 2 shows the high-level flow chart of the training and inference setup for LDM_res, while Section 4.3 holds the detailed description of LDM_res architecture.
@@
Given the incremental complexity of the tested models, in the following sections, we start by describing the core reference UNET architecture (Section 4.1), then the GAN architecture (Section 4.2), and finally LDM_res architecture (Section 4.3). The number of trainable parameters of the models is summarized in Table  1.
@@
Model Sub-net # of trainable parameters Tot # of trainable parameters

### v2

# 1]Data Science for Industry and Physics, Fondazione Bruno Kessler, via Sommarive 18, 38123, Trento (TN), Italy

@@ Abstract
Downscaling techniques are one of the most prominent applications of Deep Learning (DL) in Earth System Modeling. A robust DL downscaling model can generate high-resolution fields from coarse-scale numerical model simulations, saving the timely and resourceful applications of regional/local models. Additionally, generative DL models have the potential to provide uncertainty information, by generating ensemble-like scenario pools, a task that is computationally prohibitive for traditional numerical simulations. In this study, we apply a Latent Diffusion Model (LDM) to downscale ERA5 data over Italy up to a resolution of 2 km. The high-resolution target data consists of 2-m temperature and 10-m horizontal wind components from a dynamical downscaling performed with COSMO_CLM. Our goal is to demonstrate that recent advancements in generative modeling enable DL to deliver results comparable to those of numerical dynamical models, given the same input data, preserving the realism of fine-scale features and flow characteristics. A selection of predictors from ERA5 is used as input to the LDM, and a residual approach against a reference UNET is leveraged in applying the LDM. The performance of the generative LDM is compared with reference baselines of increasing complexity: quadratic interpolation of ERA5, a UNET, and a Generative Adversarial Network (GAN) built on the same reference UNET. Results highlight the improvements introduced by the LDM architecture and the residual approach over these baselines. The models are evaluated on a yearly test dataset, assessing the models’ performance through deterministic metrics, spatial distribution of errors, and reconstruction of frequency and power spectra distributions.
[…]
@@
High-resolution near-surface meteorological fields such as 2-m temperature and 10-m wind speed are key targets for the weather and climate scientific and operational communities. Such high-resolution information is of essential importance for a wide variety of applications (e.g. available wind potential, predicted energy consumption, etc.), across diverse timescales, from weather forecasting (nowcasting, medium-range forecasting, and seasonal predictions) to climate projections. The hunger for high-resolution data is directly linked to and justified by the information that such data hold: extreme weather events and localized phenomena can typically be described by highly resolved fields only.
@@
Downscaling is a well-known approach that allows obtaining local high-resolution data (predictands) starting from low-resolution information (predictors) by applying suitable refinement techniques. The two most traditional approaches are dynamical downscaling and statistical downscaling (Hewitson and Crane, 1996; Wilby and Wigley, 1997; Maraun and Widmann, 2018), applied alternatively depending on the final goal of each specific application.
@@
Traditionally, high-resolution fields are achieved in operational weather forecasting by performing dynamical downscaling of lower-resolution simulations. Examples of this approach are all the Local Area Models (LAMs) run in every operational center worldwide, fed with global circulation models at a low resolution, and producing high-resolution fields for a localized area, typically nationwide (e.g., Baldauf et al., 2011; Seity et al., 2011; ARPAE-SIMC, ). As for the climate community, this approach materializes, for example, in applications run within the WCRP Coordinated Regional Downscaling Experiment (CORDEX, Giorgi et al. (2009)), performing dynamical downscaling of climate projections with Regional Climate Models (RCM) going from the $\sim$100-km resolution of Global Climate Models (GCM) down to a $\sim$16-km resolution (e.g., Jacob et al., 2014, and others). The dynamical downscaling approach is well-established and provides physically and temporally consistent fields. However, it still has significant drawbacks due to the high resource demands required for its execution that limit its application, e.g. to deterministic runs instead of (or limited to small) ensemble runs.
@@
On the other hand, statistical downscaling uses coarse data from numerical simulations to infer data at high resolution by applying empirical relationships or transfer functions derived from a set of known predictors–predictands data pairs (Maraun and Widmann, 2018). Statistical downscaling methods have evolved over the years since the 1990s, with increasingly greater levels of complexity and data. Canonical examples of statistical downscaling methods are linear or multilinear regression methods (e.g., von Storch et al., 1993; Sharifi et al., 2019), analog ensemble downscaling (e.g., Sperati et al., 2024) or quantile mapping (Panofsky and Brier, 1968).
@@
In recent years, the advent of machine-learning techniques introduced many other powerful methods. These approaches have the potential to outperform classical statistical models, introducing nonlinear components in the downscaling process and learning from the provided high-resolution data. Specifically, Convolutional Neural Networks (CNNs) are particularly well suited for handling spatially distributed data and for the super-resolution task, being able to capture complex, nonlinear mappings identifying crucial features, and have already been successfully applied to weather downscaling (e.g., Baño Medina et al., 2020; Rampal et al., 2022; Höhlein et al., 2020). Building on CNN frameworks, two Deep Learning (DL) approaches are currently the most promising for improving atmospheric downscaling: Generative Adversarial Networks (GANs, Goodfellow et al. (2014, 2020)) and Diffusion Models (Sohl-Dickstein et al., 2015), which both allow for a probabilistic approach to the problem. The potential and drawbacks of these approaches are reported in the following Section 1.
@@
In this work, we focus on and test a Latent Diffusion Model (LDM) that is novel for application to the atmospheric downscaling task. Its advantages should be twofold: the diffusion approach has a much more stable training than that of GAN models, still holding the ability to generate small-scale features and the potential for ensemble production, and shows superior results in applications to image processing compared to GANs (Saharia et al., 2023; Dhariwal and Nichol, 2021). In addition, the latent approach should be seen as incremental compared to pixel-space-based diffusion approaches as it provides a cheaper solution in terms of computational costs for both inference and training (Rombach et al., 2021). This feature is intrinsically appealing to scale the downscaling task to wider (and longer) domains. Lastly, we use the high-resolution output from a numerical model dynamical downscaling simulation as our target-reference-truth. In doing so, we aim to determine whether a properly trained deep learning model can effectively emulate dynamical downscaling. If successful, this model could serve as a versatile dynamical downscaling emulator, much faster than the traditional numerical dynamical models, with a broad range of utmost significant applications.
@@ 1 Related work and Contribution
As mentioned above, currently, the most promising DL approaches for improving atmospheric downscaling are Generative Adversarial Networks (GANs) and Diffusion Models, which both are based on CNN frameworks, allow for the generation of small-scale features and for a probabilistic approach to the problem. GANs have already shown promising results in downscaling different meteorological variables, in different regions. For example, Leinonen et al. (2021) applied GANs for reconstructing high-resolution precipitation patterns from coarsened radar images; Stengel et al. (2020) demonstrated GANs potential in performing downscaling of GCMs up to 2km for solar radiation and wind; steps forward pure super-resolution applications have also been made with GANs, as shown in Harris et al. (2022) as well as in Price and Rasp (2022), where additional variables from numerical models are used as input predictors variables to produce high-resolution precipitation fields. Nevertheless, GANs still pose relevant challenges, such as model instabilities, and mode collapses during the training procedure (Arjovsky and Bottou, 2017; Mescheder et al., 2018).
@@
On the other hand, Diffusion models introduce a relatively younger approach but have already been proven very effective in weather forecasting and nowcasting applications (e.g., Leinonen et al., 2023; Li et al., 2024). Diffusion models have yet to be widely tested and evaluated on the atmospheric downscaling task, but their characteristics and capabilities are undoubtedly promising for this application as shown for example in Addison et al. (2022) and Mardani et al. (2023).
@@
Building on these encouraging results, in this work, we approach the downscaling task with a Latent diffusion model, comparing it against some standard baselines and a GAN baseline. Specifically, we re-adapted the Latent Diffusion Cast (LDCast) model (Leinonen et al., 2023), recently developed for precipitation nowcasting. LDCast has shown superior performance in the generation of highly realistic precipitation forecast ensembles, and in the representation of uncertainty, compared to traditional GAN-based methods. Our resulting model for downscaling, similar to fully convolutional models, can be trained on examples of smaller spatial domains (patches) and used at the evaluation stage on domains of arbitrary sizes, making it suitable for the generation of high-resolution data covering wider domains. As also suggested in Mardani et al. (2023), we propose the application of the diffusion model with a residual approach, relying on a standard UNET architecture for capturing the bigger scales and training the latent diffusion model to generate the residual, small scales only.
@@
Additionally, our work differs from most of the aforementioned works for the chosen pair of low-resolution and high-resolution data for the training. This choice highly influences the level of complexity that the DL downscaling model must achieve. Indeed, downscaling a coarsening of the high-resolution data (e.g., Leinonen et al., 2021; Stengel et al., 2020; Vandal et al., 2017) is a much easier task than downscaling modeled low-resolution data (e.g. short-term forecasts as in Harris et al. (2022), seasonal predictions, or climate projections) to independent high-resolution data, either coming from observations or numerical model simulation. While the first exercise falls into a purely super-resolution task, the latter includes learning potential large-scale model biases and correcting them, or detecting and generating local phenomena that cannot be resolved at the coarse resolution of the large-scale models. In our work, we focus on reanalyses products and we train our models using a set of 14 ERA5 variables as low-resolution input and high-resolution data from a dynamical downscaling of ERA5 (run with the COSMO_CLM model) as target data (2-m temperature and 10-m wind speed horizontal components). This approach is similar to that followed, for example, by Wang et al. (2021) and Mardani et al. (2023). In doing so, we intentionally force the model to learn to generate the effects of those local phenomena resolved by the dynamical numerical model, emulating its behavior.
@@ 2.1 Low- and High-resolution data
The goal of this experiment is to train a DL model to mimic a dynamical downscaling performed with a convection-permitting Regional Climate Model (RCM). The target high-resolution data consists of the hourly Italian Very High-Resolution Reanalyses produced with COSMO5.0_CLM9 (VHR-REA CCLM) by dynamically downscaling ERA5 reanalyses (Hersbach et al., 2020) from their native resolution ( 25km) up to  2.2km over Italy (Raffa et al., 2021; Adinolfi et al., 2023). Consistently with these target numerical simulations, the input low-resolution data fed to our DL model are ERA5 data.
@@ 2.2 Data alignment and preprocessing
ERA5 data have a resolution of 0.25° worldwide, which roughly corresponds to  22 km at the latitudes of the focus domain, while VHR-REA CCLM data have a native resolution of 0.02° (2.2 km). Data from both datasets were preprocessed to re-project, trim, and align the low and high-resolution fields. Specifically, the coordinate reference system (CRS) chosen for the experiment is ETRS89-LAEA Europe (Lambert Azimuthal Equal Area), also known in the EPSG Geodetic Parameter Dataset under the identifier EPSG:3035, and the experiment grids align with the European Environmental Agency Reference grid (EEA Reference grid, Peifer (2018)). ERA5 was reprojected and interpolated (with nearest-neighbor interpolation) on the EEA 16-km reference grid while VHR-REA CCLM was reprojected and interpolated on the EEA 2-km reference grid. The factor of the downscaling procedure is, therefore, 8x: over the target domain, low-resolution data consists of 72x86 16-km pixel images while high-resolution data consists of 576x672 2-km pixel images.
@@ 2.3 Experimental domain
The experiment target domain spans from 35°N to 48°N, and from 5°E to 20°E (Figure 1). This area corresponds to the region where VHR-REA CCLM data are available. The region includes a wide variety of topographically different sub-areas (mountainous areas such as the Alps and the Appennini, flat areas such as the Po Valley and coastal lines) which trigger local phenomena whose effects are challenging to identify for the downscaling models, as they are not present in the low-resolution data.
@@ 2.4 Target variables and predictors
The target variables of the study are (i) 2-m temperature and (ii) horizontal wind components at 10 m: different, dedicated models to each target variable have been trained. The input ERA5 low-resolution data are both the target variables and additional fields used as dynamical predictors, to improve models’ performance. The choice of the set of input fields was based on previous literature (e.g., Höhlein et al., 2020; Rampal et al., 2022; Harris et al., 2022) and on hardware constraints for the experiment. The selected fields used as predictor variables, for both target variables, are the following, corresponding to a total of 14 input channels to our networks:
@@
  • mean sea level pressure
@@
DEM data consists of the Copernicus Digital Elevation Model (DEM, Copernicus (2023)) interpolated from a resolution of 90 m to a resolution of 2 km. Land cover data were retrieved from the Copernicus Land Service, Global Land Cover data (GLC, Buchhorn et al. (2020)) interpolated from a resolution of 100 m to a resolution of 2 km. Given that land cover was utilized as a static variable in our analysis, we selected data from 2015: this year represents the earliest epoch available for the selected GLC dataset and falls approximately amid our experimental period of 2000-2020. Land cover class data have been converted to single-channel class masks for the DL models (totaling 16 channels). All static fields have been re-projected and aligned to the high-resolution 2-km EAA reference grid.
@@ 3 Methods
In this work, we test a deep generative Latent Diffusion Model (LDM) for the downscaling task, conditioned with low-resolution predictors and high-resolution static data. The implemented LDM is trained to predict the residual error between a previously trained reference UNET and the target variables, hence the model will be addressed as LDM_res (LDM_residual) hereafter. This residual approach has shown great performance in the application of pixel-space diffusion models (Mardani et al., 2023) and is tested here in the latent diffusion context. The underlying idea is to exploit the great ability of a relatively simple network (a UNET) to properly capture the main, bigger-scale variability of the atmospheric high-resolution data and leverage the power of the generative diffusion model to only focus on the reconstruction of the smaller-scale, locally-driven variability of the fields. Figure 2 shows the high-level flow chart of the training and inference setup for LDM_res, while Section 3.3 holds the detailed description of LDM_res architecture.
@@
Given the incremental complexity of the tested models, in the following sections, we start by describing the core reference UNET architecture (Section 3.1), then the GAN architecture (Section 3.2), and finally LDM_res architecture (Section 3.3). Table  1 shows the number of trainable parameters for each model.
@@
# of trainable parameters
Model Sub-net per subnet total